OpenStudy (anonymous):

The strength of a cable is proportional to the square of its diameter. If a 2-cm cable will support 800kg, how much will a 3-cm cable support? In Kg.

OpenStudy (anonymous):

ok so, do you know hwo many kg will 1-cm support?

OpenStudy (anonymous):

@Angelina_Vella

OpenStudy (anonymous):

are you there?

OpenStudy (lizz123):

It is just the ratios of the squares, so a 3cm cable would be 3^2/2^2 = 9/4 stronger than a 2cm cable so the weight would be 9/4 * 800 = 1800 kg

OpenStudy (lizz123):

@Angelina_Vella

OpenStudy (anonymous):

no, its 1200 kg. cuz if 1-cm= 400kg, then 3-cm=1200kg. all you have to do to find how much is in 1-cm is to cut 800kg in half since 800kg=2-cm so half of 800kg is 400kg. 1-cm=400kg, so 3-cm=1200kg. you multiply 400kg by three since you are trying to find how much 3-cm is.